OpenStudy (anonymous):

what is the difference of the following arithmetic sequence 1/2, 3/4, 1, 5/4

OpenStudy (anonymous):

First off, let's define the term, \(t_n\). \(t_n\) means the \(n\)th term in a sequence so in the case of the sequence \(\{1,5,6,10\}\), \(t_1=1\), \(t_2=5\), \(t_3=6\) and so on and so forth. In an arithmetic sequence, the common difference is a number between different terms such that: \[t_{n+1}-t_n=t_{n+2}-t_{n+1}=t_{n+3}-t_{n+2}...\] So in our case, we know: \[\eqalign{ &t_1=1/2=2/4 \\ &t_2=3/4 \\ &t_3=1=4/4 \\ &t_4=5/4 }\] So then according to our formula, \[(t_2-t_1)=(t_3-t_2)=(t_4-t_3)\] So by substituting our values: \[\left(\frac{3}{4}-\frac{2}{4}\right)=\left(\frac{4}{4}-\frac{3}{4}\right)=\left(\frac{5}{4}-\frac{4}{4}\right)=\frac{1}{4}\]